OpenStudy (anonymous):

how to solve for substitution. check your answer. y=8-x and 7=2-y

OpenStudy (anonymous):

well, let's start with the second equation 7=2-y

OpenStudy (anonymous):

what could y be?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

7 = 2 - (-5)

OpenStudy (anonymous):

so y = -5

OpenStudy (anonymous):

now you can plug in -5 for y in the first equation

OpenStudy (anonymous):

is that the answer for the first one

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

no

OpenStudy (anonymous):

-5 = 8 - x

OpenStudy (anonymous):

what is x

OpenStudy (anonymous):

-5